OpenStudy (anonymous):

a triangular window on ross camps new searay boat is two sides that measure 87cm and 64cm. The perimeter of the triangle must not exceed 291 cm. what are the possible values for the length of the third side? answer is: must be less than or equal to 140cm .

OpenStudy (anonymous):

I get 87 + 64 + x /3 >_ 291cm and get x>_431

OpenStudy (anonymous):

how do i do this right

OpenStudy (anonymous):

nvm i figured it out...

OpenStudy (anonymous):

i got 140cm btw

OpenStudy (cwrw238):

87 +64 + x <= 291

OpenStudy (anonymous):

just added the sums and didnt ever multiply by 3

OpenStudy (anonymous):

even though it doesnt make sense still.. .

OpenStudy (anonymous):

the examples in my book offer that you should divide by the # of sides

OpenStudy (cwrw238):

x <= 291-87-64 x <= 291-151 x<= 140

OpenStudy (anonymous):

thanks for helping